OpenStudy (softballgirl372015):

PLEASE HELP!!!!! If f(3)=1.8 and f(4)=-3.2, use linear interpolation to approximate f(3.6)

OpenStudy (anonymous):

i guess you are being asked to find the equation of the line between the points \((3,1.8)\) and \(( 4,-3.2)\) do you know how to do that?

OpenStudy (anonymous):

you need the slope of the line \[\frac{-3.2-1.8}{4-3}=-5\]

OpenStudy (softballgirl372015):

yes i do, but what do you do after you find the equation of the line?

OpenStudy (anonymous):

replace \(x\) by \(3.6\)

OpenStudy (softballgirl372015):

oh okay and then y = f(3.6)?

OpenStudy (softballgirl372015):

what answer did you get?

OpenStudy (anonymous):

i didn't do it, but i can. give me a second

OpenStudy (softballgirl372015):

thanks! :)

OpenStudy (anonymous):

no it isn't line is \(y=-5x+16.8\)

OpenStudy (anonymous):

then \(-5\times 3.6+16.8=-1.2\)

OpenStudy (softballgirl372015):

okay i got the same answer! thanks abillion!!!!

OpenStudy (anonymous):

yw

OpenStudy (softballgirl372015):

can you help me with another porblem?

OpenStudy (anonymous):

sure

OpenStudy (softballgirl372015):

okay here it is, is -3/2 a root of 3x^3 -5x^2-2=0

OpenStudy (anonymous):

dunno but i can use a calculator to see if \[3(\frac{-3}{2})^3-5(-\frac{3}{2})^2-2=0\]

OpenStudy (softballgirl372015):

okay

OpenStudy (softballgirl372015):

i will try tht

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3%28 \frac{-3}{2}%29^3-5%28-\frac{3}{2}%29^2-2%3D0

OpenStudy (softballgirl372015):

i used that calc and it does not equal zero. so i guess it is not a root. lol

OpenStudy (anonymous):

this one might be clearer, but it does not seem to be zero http://www.wolframalpha.com/input/?i=3%28-3%2F2%29^3-5%28-3%2F2%29^2-2

OpenStudy (softballgirl372015):

and one last question, If P is a polynomial function, and P(1)=P(4)=6, could P have a zero between 1 and 4?

OpenStudy (anonymous):

why not?

OpenStudy (softballgirl372015):

that is the answer that I got a well. you are such a help to me!!!!

OpenStudy (anonymous):

glad to help

OpenStudy (softballgirl372015):

I am not sure. lol so the answer is yes?

OpenStudy (anonymous):

yes, it could have zero between 1 and 4

OpenStudy (softballgirl372015):

okay. thank you!

OpenStudy (anonymous):

yw